# Non-Vacuous Generalization Bounds for Large Language Models

### 28 Dec 2023 v1

@@ Abstract
Modern language models can contain billions of parameters, raising the question of whether they can generalize beyond the training data or simply regurgitate their training corpora. We provide the first non-vacuous generalization bounds for pretrained large language models (LLMs), indicating that language models are capable of discovering regularities that generalize to unseen data. In particular, we derive a compression bound that is valid for the unbounded log-likelihood loss using prediction smoothing, and we extend the bound to handle subsampling, accelerating bound computation on massive datasets. To achieve the extreme level of compression required for non-vacuous generalization bounds, we devise SubLoRA, a low-dimensional non-linear parameterization. Using this approach, we find that larger models have better generalization bounds and are more compressible than smaller models.
@@ 1 Introduction
Do large language models (LLMs) merely memorize the training data, and if so, are they able to meaningfully generalize beyond their training set? This question is central to understanding LLMs as they continue to grow in capacity and are capable of memorizing and regurgitating training examples verbatim (Brown et al., 2020; Chowdhery et al., 2022; Carlini et al., 2020; 2023).
In this work, we address the question of generalization in LLMs by computing the first non-vacuous generalization bounds for language model pretraining, thereby providing a mathematical guarantee that LLMs are indeed able to generalize beyond their training data.
@@
Compared to image classification models, constructing non-trivial bounds for language models presents an additional set of challenges: (i) LLMs are trained on autoregressive token prediction, and thus token level predictions are not independent; (ii) the relevant negative log-likelihood (NLL) metric (bits per dimension) is a continuous and unbounded random variable for which previously used non-vacuous PAC-Bayes bounds are invalid; and (iii) LLMs have orders of magnitude more parameters than image classification models. To address these challenges, we derive new generalization bounds that can be applied to the unbounded bits per dimension objective, and introduce an extension of these bounds which can be computed using only a subset of the training data, substantially accelerating the bound computation for massive datasets.
Achieving the extreme level of compression required to obtain non-vacuous generalization bounds for LLMs is another challenge. To this end, we devise SubLoRA (Subspace-Enhanced Low-Rank Adaptation): a novel non-linear parameterization for LLMs that makes it possible to smoothly vary the level of compression while maintaining expressivity. SubLoRA combines low-rank adaptation (LoRA) (Hu et al., 2021), originally proposed for efficient fine-tuning, with subspace training (Li et al., 2018; Lotfi et al., 2022) to pretrain highly compressible LLMs from scratch.
Combining the above-described theoretical and practical contributions, we achieve the first non-vacuous bounds for large language models. To highlight the efficiency of our new compression technique, we compare SubLoRA to LoRA and subspace training in Figure 1 (left). This figure demonstrates that SubLoRA has an improved ability to trade off model complexity with training error. We compute bounds for both top-1 error and the bits per dimension (i.e., the average negative log-likelihood objective). The shaded region highlights where bounds become vacuous, with SubLoRA achieving non-vacuous bounds for both bits per dimension and top-1 error. In Figure 1 (right), we highlight the trade-off between model complexity and empirical risk in the generalization bounds as we vary the level of compression.
@@
Our contributions can be summarized as follows:
  1. 1.
    We design Subspace-Enhanced Lower Rank Adaptation (SubLoRA), a non-linear parameterization of the hypothesis space that enables finding compressible solutions that are sufficiently expressive to fit the training data when training from scratch. We also demonstrate the benefits of SubLoRA over LoRA and linear subspaces for next token prediction pretraining.
  2. 2.
    We derive generalization bounds for the unbounded average negative log-likelihood objective, equal to the number of bits per dimension (BPD), by incorporating prediction smoothing on the token predictions. We extend these bounds to allow for randomized subsampling of the training data in order to make the evaluation on massive datasets more tractable. Combining these contributions, we construct the first non-vacuous generalization bounds for the pretraining of large language models.
  3. 3.
    We use these bounds to address open questions about the ability of large language models to generalize beyond the training data, why models with more parameters generalize better, and the extent to which structure in the training data is coupled to the ability of LLMs to generalize.
@@ 2 Related Work
Generalization bounds. Neural networks have seen widespread adoption because of their strong performance on new unseen test samples, known as generalization. Early generalization theory literature bounded the difference in training and test error, called the generalization gap, using complexity measures like VC-dimension (Vapnik, 1991) and Rademacher complexity (Bartlett & Mendelson, 2002). These generalization bounds were vacuous for neural networks, which are often flexible enough to fit randomly labeled training data (Zhang et al., 2021). The flexibility of neural networks and its negative impact on these classical bounds calls into question why they generalize. Neural networks are so flexible that they have parameter vectors where they fit their training data and simultaneously assign incorrect labels to testing data, and they also have parameter vectors where they fit their training data and instead assign correct labels to the testing data. Why do such flexible models actually make correct test predictions in practice? Such a phenomenon can also be observed in other flexible models like Gaussian process regressors (Rasmussen & Williams, 2005), which have infinitely many parameters yet still generalize in practice.
@@
PAC-Bayes generalization theory bridges this gap by leveraging the fact that while neural networks are highly flexible and can fit random labels, they encode a preference for the correct ones (Catoni, 2007; Dziugaite & Roy, 2017). Unlike earlier generalization bounds which measured complexity merely as a function of the hypothesis class, PAC-Bayes generalization bounds reward models which have a strong prior that places its mass on parameter vectors that align with observed data. This formulation allows one to draw a parallel between generalization and compressibility (Zhou et al., 2019; Lotfi et al., 2022). By placing disproportionate prior mass on compressible parameter vectors, achieving a tight bound simply requires finding a family of models (posterior) that well fit the training data. Such compression bounds achieve the tightest guarantees to date on modern convolutional architectures and large-scale datasets, showcasing the strong inductive bias of neural networks and indicating that they can significantly compress their training sets (Lotfi et al., 2022). While PAC-Bayes has proven a very fruitful framework for devising such bounds, the insight on using a prior to bound the complexity of a given model does not require a posterior and can actually be incorporated into simpler finite hypothesis bounds.
@@
Recent generalization theory literature has expanded analysis to several relevant models—autoregressive time-series models and simple n-gram language models (McDonald et al., 2011; Bharadwaj & Hasegawa-Johnson, 2014; Vankadara et al., 2022). In contrast, we construct bounds for autoregressive transformer-based language models.
@@ 3 Background
Subspace training. We build our compression pipeline on top of techniques from several previous works. Lotfi et al. (2022) train a compressible model by parameterizing a carefully constructed low-dimensional random subspace. The weights $\theta\in\mathbb{R}^{D}$ are then defined as the sum of a random initialization $\theta_{0}$ and a projection $P\in\mathbb{R}^{D\times d}$ from a lower-dimensional subspace $w\in\mathbb{R}^{d}$: $\theta=\theta_{0}+Pw$. $P$ is constructed as the Kronecker product of random Gaussian matrices $P=(Q_{1}\otimes Q_{2})/\sqrt{D}$ for $Q_{1},Q_{2}\sim\mathcal{N}(0,1)^{\sqrt{D}\times\sqrt{d}}$, normalized so that $P^{\top}P\approx I$. The weights $w$ can then be optimized over by backpropagating through the transformation. With a learned quantization strategy—optimizing over quantized weights and the quantization levels—Lotfi et al. (2022) use arithmetic coding to encode the weights using the empirical probabilities over quantization bins.
Low Rank Adaptation (LoRA). Similarly inspired by evidence that overparametrized models have low intrinsic dimensionality (Li et al., 2018; Aghajanyan et al., 2020), Hu et al. (2021) propose LoRA as a parameter-efficient finetuning method. Given a pretrained weight matrix $W_{\text{pretrained}}\in\mathbb{R}^{a\times b}$, LoRA decomposes its total update $\Delta W$ accumulated throughout finetuning as a product of two trainable low-rank matrices $U\in\mathbb{R}^{a\times r},V\in\mathbb{R}^{r\times b}$ for $r\ll\min(a,b)$ while freezing $W_{\text{pretrained}}$. Thus $W_{\text{finetuned}}=W_{\text{pretrained}}+\Delta W=W_{\text{pretrained}}+UV$. In this work, we use LoRA for pretraining instead. In particular, we take randomly initialized neural network weights $W_{0}\in\mathbb{R}^{a\times b}$ and represent their update during pretraining as $UV$, yielding $W_{\text{pretrained}}=W_{0}+\Delta W=W_{0}+UV$. We decrease the dimensionality further by applying subspace projection to the LoRA matrices, which we describe in detail in Section 4.2.
@@ 4 Methodology
In constructing non-vacuous generalization bounds for language models, we expand and improve upon existing techniques in three ways: (1) we construct a non-linear parameterization of the compressible region of the hypothesis space which is more effective than the purely linear subspaces; (2) we construct new bounds that can handle the continuous and unbounded nature of the negative log-likelihood; (3) we make these bounds more practical to compute with LLMs by deriving a new bound which holds even when the empirical risk is evaluated only on a small subsample of the full training dataset.
@@ 4.1 Finite Hypothesis Compression Based Generalization Bounds
Given a bounded risk $R(h,x)\in[a,a+\Delta]$ and a finite hypothesis space $h\in\mathcal{H}$ for which we have a prior $P(h)$, it is straightforward to derive a generalization bound relating the empirical risk $\hat{R}(h)=\frac{1}{m}\sum_{i=1}^{m}R(h,x_{i})$ to the expected risk $R(h)=\mathbb{E}[\hat{R}(h)]$ so long as $\{x_{i}\}_{i=1}^{m}$ are sampled independently. With probability at least $1-\delta$, we have that
@@
$\log 1/P(h)\leq K(h|A)\leq C(h)\log 2+2\log C(h),$
where $C(h)$ is the compressed size of $h$ given any particular strategy for compressing $h$, where we may make use of the prior knowledge describing the architecture. Therefore, if we can find hypotheses $h$ that both have a low empirical risk and a small compressed size, then we can construct strong generalization bounds.
4.2 SubLoRA: An Efficient Non-Linear Parameterization of the Hypothesis Space
To find compressible solutions $h$ that simultaneously are expressive enough to achieve low training error, we search over a carefully designed manifold of possible parameters that live within the parameter space. In contrast to Lotfi et al. (2022), we consider a non-linear parameterization of the model weights $\theta=f(\theta_{0},w)$ given by the composition of LoRA (Hu et al., 2021) (a non-linear parameterization) and the subspace compression matrices. Given a vector of model parameters $\theta$, we break down its constituent components into the different weight matrices $W_{i}$ and associated biases $b_{i}$: $\mathrm{unflatten}(\theta)=\{(W_{i},b_{i})\}_{i\in I}$. We define a non-linear parameterization of the hypothesis space,
$\theta=\theta_{0}+\mathrm{LoRA}(Pw),$ (2)
where $\mathrm{LoRA}$ is defined by the implementation of the low rank products for the weight matrices, leaving the biases unchanged. As $Pw$ and $\theta$ are the flattened parameter vectors, $\mathrm{LoRA}(\cdot)$ is defined as the operation that unflattens the vector, applies the low rank product, and then flattens the result. For $\{(U_{i},V_{i},b_{i})\}_{i\in I}=\mathrm{unflatten}(u)$ and $\mathrm{flatten}=\mathrm{unflatten}^{-1}$, we have $\mathrm{LoRA}(u):=\mathrm{flatten}\big{(}\{(U_{i}V_{i},b_{i})\}_{i\in I}\big{)}$. Here $\theta_{0}$ is merely a random initialization of the model parameters, and $P\in\mathbb{R}^{D\times d}$ is a Kronecker product projector $P=(Q_{1}\otimes Q_{2})/\sqrt{D}$ for $Q_{1},Q_{2}\sim\mathcal{N}(0,1)^{\sqrt{D}\times d}$. We apply LoRA only over the self-attention layer and the last linear layer weight matrices, meaning that other model parameters do not differ from their initialized values. While LoRA was developed for finetuning LLMs, we find that even when training through the LoRA parameterization without changing the base parameters from their random initializations, it is possible to achieve a large fraction of the base model’s performance. Our initial exploration of LoRA for pretraining involves applying LoRA not only to attention layers but to others as well. We find that for pretraining, it is more efficient to use LoRA for both the attention layers and the last linear layer, while including other layers provides insignificant returns. In order to compress the model, we need only to represent the vector $w$ since $\theta_{0}$ and $P$ are chosen ahead of time and specified in the architecture via the random initialization.
In Figure 1 (left), we show the pareto frontier of empirical risk and the complexity penalty in the relevant generalization bound with LoRA, Subspace training, and SubLoRA. Rather than being competing methods for compression, LoRA and subspace training are complementary and exploit different structure in the parameter space to provide a family of models in the original hypothesis space that are both expressive and compressible. SubLoRA achieves a strict improvement over LoRA and subspace training, often being the deciding factor whether the bounds are vacuous or non-vacuous. In Figure 2 (left), we explore how the compressed size of the model and the empirical risk vary as a function of the subspace dimension $d$.
4.3 Accommodating the Unbounded NLL Objective Using Prediction Smoothing

### 13 Feb 2024 v2

@@ Abstract
Modern language models can contain billions of parameters, raising the question of whether they can generalize beyond the training data or simply regurgitate their training corpora. We provide the first non-vacuous generalization bounds for pretrained large language models (LLMs), indicating that language models are capable of discovering regularities that generalize to unseen data. In particular, we derive a compression bound that is valid for the unbounded log-likelihood loss using prediction smoothing, and we extend the bound to handle subsampling, making bound computation 900 times faster on massive datasets. To achieve the extreme level of compression required for non-vacuous bounds, we devise SubLoRA, a simple low-dimensional nonlinear parameterization that leads to non-vacuous generalization bounds for very large models with up to $849$ million parameters. Finally, we use our bounds to understand LLM generalization and find that larger models have better generalization bounds and are more compressible than smaller models.
@@ 1 Introduction
Do large language models (LLMs) merely memorize the training data, and if so, are they able to meaningfully generalize beyond their training set? This question is central to understanding LLMs as they continue to grow in capacity and are capable of memorizing and regurgitating training examples verbatim (Brown et al., 2020; Chowdhery et al., 2022; Carlini et al., 2020, 2023).
In this work, we address the question of generalization in LLMs by computing the first non-vacuous generalization bounds for language model pretraining on next token prediction, thereby providing a mathematical guarantee that LLMs are able to generalize beyond their training data.
@@
Compared to image classification models, constructing non-trivial bounds for language models presents additional challenges: (i) LLMs are trained on autoregressive token prediction, and thus token predictions are not independent; (ii) the relevant negative log-likelihood (NLL) metric (bits per dimension) is a continuous and unbounded random variable for which previously used non-vacuous PAC-Bayes bounds are invalid; and (iii) LLMs have orders of magnitude more parameters than image classification models. To address these challenges, we derive new generalization bounds that can be applied to the unbounded bits per dimension objective. We also introduce an extension of these bounds which can be computed using only a subset of the training data, making bound computation $900$ times faster on the OpenWebText dataset, which contains more than $9$ billion tokens.
Achieving the extreme level of compression required to obtain non-vacuous generalization bounds for LLMs is another challenge. To this end, we devise SubLoRA (Subspace-Enhanced Low-Rank Adaptation): simple nonlinear parameterization for LLMs that makes it possible to smoothly vary the level of compression while maintaining expressivity. SubLoRA combines low-rank adaptation (LoRA) (Hu et al., 2021), originally proposed for efficient fine-tuning, with subspace training (Li et al., 2018; Lotfi et al., 2022) to pretrain highly compressible LLMs from scratch.
@@
Combining the above-described theoretical and practical contributions, we achieve the first non-vacuous bounds for large language models. To highlight the efficiency of our new compression technique, we compare SubLoRA to LoRA and subspace training in Figure 1 (left). We compute two metrics that we define as follows: Top-1 Error, which is the 00-$1$ error in predicting the next token averaged over a given document; and the bits per dimension metric, which corresponds to the average negative log-likelihood per document. The shaded region highlights where bounds become vacuous, with SubLoRA achieving non-vacuous bounds for both bits per dimension and Top-1 Error. In contrast, we see that only using LoRA achieves vacuous bounds for Top-1 Error and only using subspace achieves a high value of empirical BPD. Despite the simplicity of SubLoRA, it has an improved ability to trade-off model complexity with training error. In Figure 1 (right), we highlight the trade-off between model complexity and empirical risk in the generalization bounds as we vary the level of compression.
We summarize our contributions as follows:
  • Novel bounds for the unbounded negative log-likelihood objective: we introduce novel bounds specifically tailored to account for the unbounded continuous bits-per-dimension loss, commonly used to evaluate LLMs for next-token prediction.
  • Subsampling bounds for practical bound evaluation: To make the evaluation of the bounds practical on LLMs with massive datasets, we derive subsampling-based bounds that allow for efficient evaluation. In practice, the evaluation of the bound takes 45 minutes on a single GPU instead of 3 days on 8 GPUs in parallel for the OpenWebText dataset.
  • A simple yet powerful nonlinear subspace compression for LLMs: as we show in Figure 1, using LoRA alone to compress the discrepancy between the random initialization and a learned model leads to vacuous bounds for the top-1 error. At the same time, linear subspace training alone does not unlock the full compression potential of LLMs compared to a nonlinear compression scheme. We show that a combination of these two approaches, while simple, yields a strong nonlinear compression of the model, which leads to the best generalization bounds for LLMs.
  • Non-vacuous generalization bounds for models with nearly a billion parameters: our work does not only introduce the first non-vacuous generalization bounds for LLMs, but it also extends these bounds to models with over $800$ million parameters, demonstrating the scalability of our compression technique.
  • Improved understanding of generalization in LLMs: as we increase the size of models, we find that they are more compressible and achieve better bounds, therefore disproving the claim that larger LLMs are simply better at regurgitating their training data.
The significance of these contributions lies in the ability to offer mathematical proof that large language model are, in fact, powerful knowledge compressors and are capable of generalization beyond their training samples, especially as their scale increases. To the best of our knowledge, our work is the first to show that generalization bounds improve with more parameters on models of practical sizes, in line with the empirical benefits of large models.
@@ 2 Related Work
Generalization bounds. Neural networks have seen widespread adoption because of their strong performance on new unseen test samples, known as generalization. Early generalization theory literature bounded the difference in training and test error, called the generalization gap, using complexity measures like VC-dimension (Vapnik, 1991) and Rademacher complexity (Bartlett and Mendelson, 2002). These generalization bounds were vacuous for neural networks, which are often flexible enough to fit randomly labeled training data (Zhang et al., 2021). The flexibility of neural networks and its negative impact on these classical bounds calls into question why they generalize. Neural networks are so flexible that they have parameter vectors where they fit their training data and simultaneously assign incorrect labels to testing data, and they also have parameter vectors where they fit their training data and instead assign correct labels to the testing data. Why do such flexible models actually make correct test predictions in practice?
@@
PAC-Bayes generalization theory bridges this gap by leveraging the fact that while neural networks are highly flexible and can fit random labels, they encode a preference for the correct ones (Catoni, 2007; Dziugaite and Roy, 2017; Arora et al., 2018). Unlike earlier generalization bounds which measured complexity merely as a function of the hypothesis class, PAC-Bayes generalization bounds reward models which have a strong prior that places its mass on parameter vectors that align with observed data. This formulation allows one to draw a parallel between generalization and compressibility (Zhou et al., 2019; Lotfi et al., 2022). By placing disproportionate prior mass on compressible parameter vectors, achieving a tight bound simply requires finding a family of models (posterior) that well fit the training data. Such compression bounds achieve the tightest guarantees to date on modern convolutional architectures and large-scale datasets, showcasing the strong inductive bias of neural networks and indicating that they can significantly compress their training sets (Lotfi et al., 2022). While PAC-Bayes has proven a very fruitful framework for devising such bounds, the insight on using a prior to bound the complexity of a given model does not require a posterior and can actually be incorporated into simpler finite hypothesis bounds.
@@
Recent generalization theory literature has expanded analysis to several relevant models—autoregressive time-series models and simple n-gram language models (McDonald et al., 2011; Bharadwaj and Hasegawa-Johnson, 2014; Vankadara et al., 2022). In contrast, we construct bounds for autoregressive transformer-based language models.
Existing bounds for unbounded objectives. A number of works have explored techniques for generating generalization bounds on unbounded objective functions more generally, but these approaches are not practical for application to LLMs. A well established strategy relevant for e.g. linear regression with Gaussian errors is to bound the tails of the objective as subgaussian random variables, and then generalization bounds can be constructed for subgaussians more generally (Alquier et al., 2016; Germain et al., 2016). Other kinds of known tail behavior have also been exploited (Holland, 2019; Kuzborskij and Szepesvári, 2019). For the NLL of a language model, there is no clear analogous tail behavior, so we must take a different approach.
Haddouche et al. (2021) devise an approach for general unbounded objectives by constructing a hypothesis dependent bound on the objective, even if the objective is unbounded more generally. If the risk can be bounded $\sup_{x}R(h,x)\leq Q(h)$ for a function $Q(h)$, then PAC-Bayes bounds can be constructed using $Q(h)$ even if $\sup_{h}Q(h)=\infty$. However, even though $Q(h)$ is finite for LLMs as there are only a finite number of inputs, $Q$ grows exponentially for NLL with the number of layers in the network and is closely related with the Lipschitz constant. For large models like LLMs, this value is far too large to be useful in constructing bounds.
@@ 3 Background
Subspace training. Lotfi et al. (2022) train a compressible model by parameterizing a carefully constructed low-dimensional random subspace. The weights $\theta\in\mathbb{R}^{D}$ are then defined as the sum of a random initialization $\theta_{0}$ and a projection $P\in\mathbb{R}^{D\times d}$ from a lower-dimensional subspace $w\in\mathbb{R}^{d}$: $\theta=\theta_{0}+Pw$. $P$ is constructed as the Kronecker product of random Gaussian matrices $P=(Q_{1}\otimes Q_{2})/\sqrt{D}$ for $Q_{1},Q_{2}\sim\mathcal{N}(0,1)^{\sqrt{D}\times\sqrt{d}}$, normalized so that $P^{\top}P\approx I$. The weights $w$ can then be optimized over by backpropagating through the transformation. With a learned quantization strategy—optimizing over quantized weights and the quantization levels—Lotfi et al. (2022) use arithmetic coding to encode the weights using the empirical probabilities over quantization bins.
Low Rank Adaptation (LoRA). Similarly inspired by evidence that overparametrized models have low intrinsic dimensionality (Li et al., 2018; Aghajanyan et al., 2020), Hu et al. (2021) propose LoRA as a parameter-efficient finetuning method. Given a pretrained weight matrix $W_{\text{pretrained}}\in\mathbb{R}^{a\times b}$, LoRA decomposes its total update $\Delta W$ accumulated throughout finetuning as a product of two trainable low-rank matrices $U\in\mathbb{R}^{a\times r},V\in\mathbb{R}^{r\times b}$ for $r\ll\min(a,b)$ while freezing $W_{\text{pretrained}}$. Thus $W_{\text{finetuned}}=W_{\text{pretrained}}+\Delta W=W_{\text{pretrained}}+UV$. In this work, we use LoRA for pretraining instead. In particular, we take randomly initialized neural network weights $W_{0}\in\mathbb{R}^{a\times b}$ and represent their update during pretraining as $UV$, yielding $W_{\text{pretrained}}=W_{0}+\Delta W=W_{0}+UV$. We decrease the dimensionality further by applying subspace projection to the LoRA matrices, which we describe in detail in Section 5.
@@ 4 Methodology
In constructing non-vacuous generalization bounds for LLMs, we expand and improve upon existing techniques in three ways: (1) we construct a simple and effective nonlinear parameterization which is more effective and scalable than purely linear subspaces; (2) we construct new bounds that can handle the continuous and unbounded nature of the negative log-likelihood; (3) we make these bounds more practical to compute with LLMs by deriving a new bound which holds even when the empirical risk is evaluated only on a small subsample of the full training dataset.
@@ 4.1 Finite Hypothesis Compression Based Generalization Bounds
Given a bounded risk $R(h,x)\in[a,a+\Delta]$ and a finite hypothesis space $h\in\mathcal{H}$ for which we have a prior $P(h)$, it is straightforward to derive a generalization bound relating the empirical risk $\hat{R}(h)=\frac{1}{m}\sum_{i=1}^{m}R(h,X_{i})$ to the expected risk $R(h)=\mathbb{E}[\hat{R}(h)]$ so long as $\{X_{i}\}_{i=1}^{m}$ are sampled independently. With probability at least $1-\delta$, we have
@@
$\log 1/P(h)\leq K(h|A)\log 2\leq C(h)\log 2+2\log C(h),$
where $C(h)$ is the compressed size of $h$ given any particular strategy for compressing $h$ and we may make use of the prior knowledge describing the architecture. Therefore, if we can find hypotheses $h$ that both have a low empirical risk and a small compressed size, then we can construct strong generalization bounds.
4.2 Enabling the Independence Assumption for Generalization
Bounds on Text Data
Using Equation 1 requires that $X_{i}$ in the sum $\hat{R}(h)=\frac{1}{m}\sum_{i=1}^{m}R(h,X_{i})$ are drawn independently. Thus, we must be careful in the construction and interpretation of our bounds so that this constraint is satisfied. Instead of considering bounds at the level of tokens, which are correlated, we instead define $X_{i}$ to be an entire document sampled uniformly from the data generating process from which the corpus was sampled. In other words, we break our dataset into its constituent documents and sample uniformly documents uniformly from it, where each $X_{i}$ represents an entire document. We define the risk on a given document as the negative log-likelihood of the entire document divided by its length, according to the autoregressive model.
It is also possible to choose $X_{i}$ to be a context chunk, i.e., a sequence of length equal to the context length, as is commonly used in the training of models since a document may be larger than the maximum transformer context length. In such cases, the sequences are no longer independent samples from the data generating process. It is possible to construct valid bounds on these sequences which respect the independence assumption. However, in doing so we must shift the interpretation of the bounds from being over the randomness in sampling from the data generating process to the randomness in sampling sequences that can be constructed from a fixed and finite dataset formed by concatenating the documents together. We explore these alternate sequence-level bounds in Appendix B. However, we believe that the document-level bounds provide a more meaningful and significant statement about generalization.
4.3 Accommodating the Unbounded Negative Log-Likelihood
Objective Using Prediction Smoothing